Think of a number. Add 10. Multiply by 4. Subtract 18. Divide by 2. Write down two additional steps which will always bring you back to the number you started with.
Answer by checkley77(12844)

6+10=16*4=64-18=46/2=23[-5]=18/[3]=6 back home.